Andrea's savings account has a balance of $504. After 4 years, what will the amount of interest be at 5% compounded quarterly?
Answer by stanbon(48558)

Andrea's savings account has a balance of $504. After 4 years, what will the amount of interest be at 5% compounded quarterly?
-----------------
A(t) = P(1+(r/n))^(nt)
504 = P(1+(0.05/4))^(4*4)
504 = P(1.2199)
504 = P + P(0.2199)
-------------
Interest = 0.2199 or 21.99%
==============================
Cheers,
Stan H.